OpenStudy (anonymous):

3x - 2t = 4 2x -5 = -t

OpenStudy (whovianchick):

3x - 2t = 4 2x -5 = -t 3x - 2t = 4 3x-2t+2t=4+2t 3x=2t+4 x=(2/3)t+(4/3) 2x-5=-t 2[(2/3)t+(4/3)]-5=-t (4/3)t+(8/3)-5=-t (8/3)-5=-(7/3)t -7/3=-7/3t t=1 2x -5 = -t 2x -5 = (1) 2x-5+5=1+5 2x=6 x=3 t=1 x=3

OpenStudy (countryboy15):

hi @whovianchick

OpenStudy (anonymous):

Thanks @whovianchick

OpenStudy (whovianchick):

No prob, sorry it was super long I just like to show all my work so you can see the process, not just the answer

jaynator495 (jaynator495):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @whovianchick 3x - 2t = 4 2x -5 = -t 3x - 2t = 4 3x-2t+2t=4+2t 3x=2t+4 x=(2/3)t+(4/3) 2x-5=-t 2[(2/3)t+(4/3)]-5=-t (4/3)t+(8/3)-5=-t (8/3)-5=-(7/3)t -7/3=-7/3t t=1 2x -5 = -t 2x -5 = (1) 2x-5+5=1+5 2x=6 x=3 t=1 x=3 \(\color{#0cbb34}{\text{End of Quote}}\) As this was a good answer, it also was !!THE!! answer, in the future, please guide them through the process, and don't flat out give them the answer, this will actually teach them how to do a similar question later on... Thank you ^.^